# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v4

@@ A Level Total Order (LTO) on a set $S$ can be defined as follows:
    $u$ is less than $v$ (denoted as $u<v$), $u$ is greater than $v$ (denoted as $u>v$), or $u$ is equal to $v$ (denoted as $u=v$).
  • Leveled Structure: The elements of $S$ are partitioned into a leveling of ”containers” or ”levels,” $C_{1},C_{2},,,,,C_{n}$, with each container itself forming a subset of $S$ that maintains a total order. The containers are sequentially ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
  • Comparability within Containers: For any two elements $u,v\in Ci$, their order can be determined based on a specific criterion, such as a designated metric or property relevant to the application.
@@
Decreasing Sequence Property Lemma If a node in an oriented graph $G$ does not have the decreasing sequence property, then its out-degree at least doubles in $G^{2}$.
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$, and those that are not neighbors of $x$. In the proofs of Lemma 4.2(Minimum Out-Degree 3 with Neighbor Out-Degree 0), and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor induced subgraph (of $v_{0}$) and in the graph with the neighbor induced subgraph removed. We will now give terms to these concepts.
@@
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
@@
,where $N^{+}(z)$ is the set of out-neighbors of $z$. The number of interior neighbors is the interior degree of $y$ with respect to $x$.
Now that we have introduced containers, we can talk arcs where both nodes are inside the containers, these are interior neighbors. Interior neighbors will be important to our study on Seymour’s Second Neighborhood Conjecture and our search for a counterexample.
In particular we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1 and we see that when all the neighbors in that subgraph had outdegree of 1, $v_{0}$’s degree did not double.
@@
, where $N^{+}(y)$ is the set of out-neighbors of $y$, and $N^{++}(x)$ is the set of second neighbors of $x$. The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
@@
As an exercise to visualize interior and exterior neighbors, consider example 5.2. Exterior arcs are the double arrows between the boxes. Interior arcs are not drawn but can be mentally placed either between any two nodes in the red or blue box.
@@
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. In order to do so though, we need to consider the option of back-arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to the minimum degree node. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture where they may force us to double count nodes.
Having defined our arcs, we can talk about the decreasing sequence property. This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decreases as we get further from the minimum degree node. Exterior degrees, though, are a part of a node’s degree equation. So there’s another part of it. As we move further from the minimum degree node, the size of the interior degrees increases. This phenomenon reflects a yin and yang relationship between the interior and exterior degrees.
In all of this, the minimum degree node is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s outdegrees.
With this definition of interior and exterior nodes now introduced, we are now ready to now introduce the Graph Level Order (GLOVER) data structure.
@@
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Leveled Structure: The elements of $S$ are partitioned into levels of ”containers” or ”rankings,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
@@ A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Given an oriented graph $G$ and a minimum degree node $v_{0}$, where the nodes of $G$ are partitioned into containers $C_{1},C_{2},\dots,C_{k}$ based on their distance from $v_{0}$ and assuming no back arcs, and the interior arcs in each container and exterior arcs between each container forms a Graph Level Order (GLOVER) under the distance total order.
@@
To establish that the structure $\{C_{1},C_{2},\dots,C_{k}\}$ satisfies the GLOVER properties, we verify each condition in the GLOVER definition:
  • Total Order on Elements within $G$: - By partitioning nodes based on their distance from $v_{0}$, we establish a total order on the elements of $G$: for any two nodes $u$ and $v$, their order is determined by their respective containers. Nodes in closer containers (e.g., $C_{i}$) are considered ”less than” nodes in further containers (e.g., $C_{j}$ with $i<j$).
    - Within each container, nodes may be ordered based on additional criteria (such as their individual distances to $v_{0}$ if there are multiple paths). Thus, we achieve a total order among all nodes in $G$.
  • Leveled Structure with Containers: - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$. The elements within each container are in a total order relative to each other, and the containers themselves are ordered by distance from $v_{0}$, establishing a multi-level ranked structure.
  • Non-triviality Requirement: - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node. This fulfills the non-triviality requirement, as we avoid having all singleton containers.
  • Comparability within Containers: - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$ (for example, their shortest-path distance if there are multiple paths). Thus, comparability is ensured within containers.
  • Consistency of Comparisons Across Containers: - Since the graph is oriented and there are no back arcs (i.e., no edges from nodes in $C_{j}$ to nodes in $C_{i}$ for $j>i$), the ordering between containers is respected. This guarantees that if $C_{i}$ is ”less than” $C_{j}$, then every element in $C_{i}$ is ”less than” every element in $C_{j}$, ensuring consistency across containers.
  • Interior and Exterior Neighbors: Nodes $u$ and $v$ of the same distance from the chosen minimum degree node $v_{0}$ would be in the same container and thus be interior neighbors. If these nodes were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance different by 1.
    Each criterion for an GLOVER is satisfied by the structure defined by partitioning nodes based on distance from $v_{0}$. Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a Ranked Total Order under the distance total order, completing the proof.
@@
Next, we will show the power of this data structure and why is introduced in this paper.
@@
GLOVER Supports Induction: In any oriented graph, the minimum degree node $v_{0}$ within a Glover structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
@@
Let $v_{0}$ be the node with the minimum out-degree in the oriented graph $G$. By definition, for any node $v\in G$, $d^{+}(v_{0})\leq d^{+}(v)$, where $d^{+}(v)$ is the out-degree of node $v$. This ordering provides a baseline for comparison because all other nodes are either equal to or have greater out-degree than $v_{0}$.
The analysis begins at $v_{0}$, which serves as the foundation of the GLOVER structure. All outgoing arcs from $v_{0}$ lead to nodes at the next level (or ”container”) $C_{1}$. By the properties of the GLOVER structure (e.g., no back arcs), the relationships between nodes are strictly directed, ensuring no circular dependencies.
Assume that for any node $v$ in a container $C_{i}$, the following properties hold:
  • $d^{+}(v)\geq d^{+}(v_{0})$.
  • All properties derived from $v_{0}$ apply recursively to nodes in $C_{i}$.
Consider any node $u\in C_{i+1}$ that is connected to a node $v\in C_{i}$. By the inductive hypothesis, $v$ satisfies all conditions derived from $v_{0}$. Since $u$ receives its inbound arcs from $v$, and $d^{+}(v)\geq d^{+}(v_{0})$, the relationships in $C_{i+1}$ are consistent with the properties of the Glover structure. Furthermore, no node $u$ in $C_{i+1}$ can violate the inductive assumptions because $v_{0}$ ensures a well-ordering, propagating constraints downward.
Starting from $v_{0}$, the well-ordering propagates through all containers $C_{i}$, ensuring that:
  • The out-degree of any node is anchored to $v_{0}$.
  • There is no upward deviation in degree that contradicts the structure’s constraints.
Thus, the GLOVER structure supports induction by leveraging the minimum degree node $v_{0}$ as an anchor, ensuring consistency throughout the graph. ∎
@@
Figure 6: This example illustrates back arcs in a graph. The two blue arcs from $x3$ to $y3$ and $y3$ to $z3$ represent exterior neighbors indicate forward directing arcs. Additionally, a red arc points backwards from $z5$ to $x1$. Although the containers are not outlined, they are implied by the difference in variables for each group.
@@
Interior Degree Greater than 1 In any oriented graph $G$, represented by a GLOVER with $v_{0}$ as a minimum out-degree node satisfying the decreasing sequence property, In an attempt to prevent a degree doubling node, every node $x$ in the neighboring container $N^{+}(v_{0})$ has an interior degree of at least 1. That is, for $x\in N^{+}(v_{0})$,
@@
By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$
where $N^{+}(x)$ denotes the set of first neighbors (out-neighbors) of $x$, and $N^{++}(x)$ denotes the set of second neighbors (out-neighbors of $N^{+}(x)$).
For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $y$, and then from $y$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $y\in N^{+}(v_{0})$.
If $y$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $y$’s neighbors, which are $v_{0}$’s second neighbors. If $y$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $y$’s first neighbors.
To avoid this doubling, $y$ must connect to at least one interior arc. This means that the interior degree of $y$ with respect to $v_{0}$ must satisfy
$|int(v_{0},y)|\geq 1$
This argument holds for any neighbor $y\in N^{+}(v_{0})$, ensuring that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
Thus, for all $x\in N^{+}(v_{0})$,$|int(v_{0},x)|\geq 1$, completing the proof. ∎
Lemma 5.4 (Interior Degree Greater than 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling. In particular, we see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree to double.
@@
Exterior Less than Degree In any oriented graph $G$, represented by a GLOVER In an attempt to prevent a degree doubling node, if a node $x_{i}$ is in container $C_{i}$ has the decreasing sequence property, In an attempt to prevent a degree doubling node, then for all $y\in N^{+}(x)\subseteq C_{i+1}$, $|ext(x,y)|<d^{+}(x)$.
@@
Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|ext(x,y)|\geq d^{+}(x)$. By definition, $ext(x,y)$ consists of the set of out-neighbors of $y$, specifically:
$ext(x,y)=\{z\in N^{++}(x)\mid z\text{ is an out-neighbor of }y\}.$
Thus, $|ext(x,y)|\subseteq N^{++}(x)$, where $N^{++}(x)$ denotes the set of second out-neighbors of $x$. If $|ext(x,y)|\geq d^{+}(x)$, then:
@@
which suggests that the out-degree of $x$ would at least double in the graph $G^{2}$ (the second power of the graph, where edges represent paths of length 2). This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property, where the degree of nodes cannot increase as we move through the containers.
Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|ext(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
Lemma 5.5 is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree Greater than 1), with the notable exception that Lemma LABEL:extbd1 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
The results presented in 5.4(Interior Degree Greater then 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1. There, he demonstrated that a minimum degree vertex with no cycles in its induced graph has a degree that effectively doubles.
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we’re developing.
@@
[…]
Then, for all $z\in ext(x,y)$, we have $|ext(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
@@
Now, assume the conditions of the lemma hold: 1. The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increases significantly in $G[N^{+}(x)]$. 2. For all $z\in ext(x,y)$, $ext(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
Now, suppose for contradiction that there exists $z\in ext(x,y)$ such that $|ext(y,z)|\geq|ext(x,y)|$. This would mean that the exterior degree of $y$ is at least as large as the exterior degree of $x$. However, this contradicts the assumptions about the DSP of $y$, which requires $y$ to have fewer second neighbors than first neighbors.
@@
Thus, the assumption $|ext(y,z)|\ \geq|\text{ext}(x,y)|$ leads to a contradiction,
confirming that $|\text{ext}(y,z)|<|\text{ext}(x,y)|$ for all $z\in\text{ext}(x,y)$.
Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$. ∎
The first thing to notice in this proof are the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions has an important purpose.
First we will talk about the lack of back arcs. We have already talked about how back arcs can be a hindrance to our algorithm overall. This is one of the main points where they can be a hindrance. In particular there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist. We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
Secondly, we assumed that a node’s interior degree doubled. This is similar to the notion of solving the containers separately, then joining them together again to solve the main problem. We are assuming that this not $y$ was in another container and its interior degree doubled in that container. In doing so, $y$ is now in a better position to have its overall degree double, needing only to have its exterior degree double.

### v6

@@ A Level Total Order (LTO) on a set $S$ can be defined as follows:
    • $u$ is less than $v$ (denoted as $u<v$),
    • $u$ is greater than $v$ (denoted as $u>v$), or
    • $u$ is equal to $v$ (denoted as $u=v$).
  • Leveled Structure: The elements of $S$ are partitioned into a leveling of containers. $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$ Each container itself forms a subset of $S$. This maintains a total order. The containers are sequentially ordered relative to each other. This establishes a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
  • Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$. This avoids trivial total orderings where all elements are isolated in singleton containers.
  • Comparability within Containers: For any two elements $u,v\in C_{i}$, their order can be determined based on a specific criterion.
  • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves.
@@
Decreasing Sequence Property Lemma Suppose a node in an oriented graph $G$ does not have the decreasing sequence property. Then this node will have its out-degree at least double in $G^{2}$.
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$. In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed. We will now give terms to these concepts.
@@
Let $x$ be a node with the decreasing sequence property and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
@@
The number of interior neighbors is the interior degree of $y$ with respect to $x$.
Now that we have introduced containers, we can talk about arcs where both nodes are inside the containers; these are interior neighbors. Interior neighbors will be important to our study on SSNC and our search for a counterexample.
In particular, we saw things about interior neighbors in 4 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
@@
The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
@@
As an exercise to visualize interior and exterior neighbors, consider example 5.2. Exterior arcs are the double arrows between the boxes. Interior arcs are not drawn but can be mentally placed either between any two nodes in the container boxes in example 5.2 or example 5.1.
@@
Let $v_{0}$ be a minimum degree node. Suppose that $x$ is a node in the component $C_{i}$. A back arc is defined as an arc $(x,y)$ such that $y\in N^{+}(x)$ and $y\in C_{j}$, where $j<i$.
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. In order to do so, though, we need to consider the option of back-arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to the minimum degree node. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes.
Having defined our arcs, we can talk about the decreasing sequence property. This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from the minimum degree node. Exterior degrees, though, are a part of a node’s degree equation. There’s another part of it though. As we move further from the minimum degree node, the size of the interior degrees increases.
This phenomenon reflects a yin and yang relationship between the interior and exterior degrees. In all of this, the minimum degree node is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees. With this definition of interior and exterior nodes now introduced, we are now ready to introduce the Graph Level Order (GLOVER) data structure.
@@
A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
Leveled Structure: The elements of $S$ are partitioned into levels of containers $C_{1},C_{2},\ldots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$.
@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
Assume we have an oriented graph $G$. Assuming that there are no back arcs in $G$, this oriented graph can be represented by a GLOVER.
@@
To establish a Graph Level Order, we first need a minimum degree node. We can obtain this by taking the minimum value on the degree vector of nodes. We will call that in the node $v_{0}$.
Now that we know the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$.
Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes whose distance from the node $v_{0}$ is equal. These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
  • Total Order on Elements within $G$ Because nodes are partitioned based on their distance from $v_{0}$, we have a total order on the elements of $G$. For any two nodes $u$ and $v$, this total order is determined by their containers.
  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$.
  • Non-triviality Requirement - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node.
  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$.
  • Consistency of Comparisons Across Containers - Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
  • Interior and Exterior Neighbors:
    • Nodes $u$ and $v$ of the same distance from $v_{0}$ would be in the same container and thus be interior neighbors.
    • If the nodes $u$ and $v$ were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance differs by 1.
    Each criterion for a GLOVER is satisfied by partitioning nodes based on distance from $v_{0}$. Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a GLOVER under the distance total order, completing the proof.
@@
Next, we will show one of the other benefits of this data structure in mathematical induction via well ordering.
@@
GLOVER Supports Induction
In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
@@
We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering.
A set is considered well-ordered if every non-empty subset of that set has a least element.
A discrete total order is a total order where each element has a distinct ”next” element in the order.
Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order. This fulfills the condition of being well-ordered.
Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set.
@@
Figure 6: This example illustrates back arcs in a graph. The two blue arcs from $x_{3}$ to $y_{3}$ and $y_{3}$ to $z_{3}$ represent exterior neighbors indicating forward-directing arcs. Additionally, a red arc points backwards from $z_{5}$ to $x_{1}$. Although the containers are not outlined, they are implied by the difference in variables for each group.
@@
Interior Degree at least 1
Suppose we have an oriented graph $G$, represented by a GLOVER. Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property. In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $N^{+}(v_{0})$ must have an interior degree of at least 1. That is, for $x\in N^{+}(v_{0})$,
@@
By definition, a node satisfies the decreasing sequence property if
$|N^{+}(x)|>N^{++}(x)|$
For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $x$, and then from $x$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $x\in N^{+}(v_{0})$.
If $x$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $x$’s neighbors, which are $v_{0}$’s second neighbors. If $x$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $x$’s first neighbors.
To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
$|int(v_{0},x)|\geq 1$
This argument holds for any neighbor $x\in N^{+}(v_{0})$. This ensures that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
Thus, for all $x\in N^{+}(v_{0})$, $|int(v_{0},x)|\geq 1$.
Lemma 5.4 (Interior Degree at least 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling. In particular, we see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree to double.
@@
Exterior Less than Degree
Suppose we have an oriented graph $G$, represented by a GLOVER. In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i}),|\text{ext}(x,y)|<d^{+}(x)$.
@@
Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$.
By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$
Thus, $|\text{ext}(x,y)|\subseteq N^{++}(x)$, then:
@@
suggesting that the degree of outwardness of $x$ would at least double on graph $G^{2}$. This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property.
Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$.
Lemma 5.5 (Exterior Less than Degree) is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree at least 1), with the notable exception that Lemma 5.5 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
The results presented in 5.4 (Interior Degree At Least 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1. There, he demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we are developing.
@@
Weak Decreasing Exteriors Lemma
Suppose that $G$ is an oriented graph represented by a GLOVER, and that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
Suppose that:
  1. 1.
    The interior degree of $y$ with respect to $x$, $|int(x,y)|$, doubles in the subgraph $G[N^{+}(x)]^{2}$.
  2. 2.
    For all $z\in ext(x,y)$, we have $ext(y,z)\cap N^{+}(x)=\emptyset$, i.e., no back arcs from the second neighbors of $x$ to the first neighbors of $x$ exist.
Then, for all $z\in\text{ext}(x,y)$, we have $|\text{ext}(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
@@
Now, assume the conditions of the lemma hold:
  1. 1.
    The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increase significantly in $G[N^{+}(x)]$.
  2. 2.
    For all $z\in\text{ext}(x,y)$, $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
Now, suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$.
The node $y$ has a degree that can be broken down into two components, its interior degrees and its exterior degrees. By the assumption that $y$’s interior degree doubles, if $y$’s exterior degree were to double, it would lead to a contradiction.
If there exists a node $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$, then $y$’s out-degree will be impacted by $\text{ext}(x,y)$ and $\text{ext}(y,z)$.
$\begin{split}|\text{ext}(y,z)|&\geq|\text{ext}(x,y)|\\ |\text{ext}(y,z)|+|ext(x,y)|&\geq|\text{ext}(x,y)|+|\text{ext}(x,y)|\\ |\text{ext}(y,z)|+|\text{ext}(x,y)|&\geq 2*|\text{ext}(x,y)|\end{split}$ (1)
This would mean that the exterior degree of $y$ is at least as large as the exterior degree of $x$. However, this contradicts the assumptions about the DSP of $y$, which requires $y$ to have fewer second neighbors than first neighbors.
@@
Thus, the assumption $|\text{ext}(y,z)|\allowbreak\geq\allowbreak|\text{ext}(x,y)|$ leads to a contradiction, confirming that $|\text{ext}(y,z)|\allowbreak<\allowbreak|\text{ext}(x,y)|$ for all $z\allowbreak\in\allowbreak\text{ext}(x,y)$.
Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$.
The first thing to notice in this proof is the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions are important.
First, we will talk about the lack of back arcs. We have already talked about how back arcs can be a hindrance to our algorithm overall. This is one of the main points where they can be a hindrance. In particular, there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist. We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
Secondly, we assumed that the interior degree of a node doubled. This is similar to the notion of solving the containers separately, then joining them together again to solve the main problem. We are assuming that this node $y$ was in another container and its interior degree doubled in that container. In doing so, $y$ is now in a better position to have its overall degree double, only needing to have its exterior degree double.
The more popular representations of graphs are adjacency matrices, arraylists and edge lists. These suffer from the same limitations of graphs as mentioned above, They scatter the nodes along the two-dimensional plane without attempts to group into an ordering. The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. In order to represent graphs in a GLOVER we would use any markup language that can encode graphs. There needs to be three types of data- containers, nodes, and arcs. The containers will be given ids, the nodes will then be assigned to their unique containers, and arcs will be determined to be interior, exterior, or back, depending on the containers of the endpoints of the arcs.
This determination of the number of containers and the assignment of nodes to containers does not need to be declared beforehand. Instead, just as we have been talking about the distance of a node from a minimum degree node $v_{0}$ as the determination of container assignment, we would have the same or a similar function to assign nodes to containers.
With this container representation of the SSNC, by way of the GLOVER data structure, we can begin to visualize elements of these lemmas. For example, the lemma 5.6 (Weak Decreasing Exteriors Lemma) can be visualized by example 5.5. This shows a series of ordered containers. Each container has fewer nodes. The fewer nodes implies that the container circle can be drawn with a smaller radius, giving light to the decrease in size.